please help me in finding the integral of
     ∫  1/(x√x4-1) dx.

Grade:12

1 Answers

Aditya Gupta
5 years ago
Multiply and divide by x^3
So we have
x^3/x^4(√(x^4+1))
Now put x^4=y
Now to calculate dy/y√(y+1), simply put y+1=z^2 and then use partial fractions.